ANALYSIS OF ROTATIONAL STRUCTURE IN THE HIGH-RESOLUTION INFRARED SPECTRUM OF cis-1,3,5-HEXATRIENE

\underline{NORMAN C. CRAIG}, and MATTHEW C. LEYDEN, Department of Chemistry and Biochemistry, Oberlin College, Oberlin, OH 44074; TONY MASIELLO, Environmental and Molecular Sciences Laboratory, Pacific Northwest National Laboratory, Richland, WA 99352.

The high-resolution infrared spectrum has been recorded for two C-type bands of cis-1,3,5-hexatriene. The resolution (0.0013 cm^{−1}) and the Doppler width (0.0012 cm^{−1} at 900 cm^{−1}) are barely adequate for observing detailed rotational structure of this heavy molecule in a spectrum recorded at room temperature. An additional complication is the extensive hot band structure arising from excited states of the low frequency C-C torsional modes. A preliminary analysis of rotational structure yielded ^{R}R_{K} and ^{P} P_{K} series in each of the two bands, centered at 908 cm^{−1} and 586 cm^{−1}. However, ground state combination differences (GSCDs) failed to fit a rotational Hamiltonian. Subsequent microwave spectroscopic measurements gave ground state rotational constants. Reliable GSCDs computed from the ground state rotational constants led to revisions in assignments of some of the sub-band series in the infrared spectrum and to a convincing assignment.